OpenStudy (anonymous):

A manufacturer determines that the number of drills it can sell is given by the formula D= -3p^2 + 180p - 285, where p is the price of the drills in dollars. a.) At what price will the manufacturer sell the maximum number of drills? b.) What is the maximum number of drills that can be sold?

OpenStudy (anonymous):

For (a), you should set the derivative of D with respect to p, to zero. You get -6p + 180 = 0. So, p = 180/6 = 30.

OpenStudy (anonymous):

For (b), plug that value of p = 30 into the equation for D and you will get a number.